AB INIT IO POTENTIAL ENERGY SURFACES FOR THE I(^{2}P_{3/2})+ O_{2}(a^{1}Δ_{g})↔I(^{2}P_{1/2})+ O_{2}(X^{3}Σ^{−}_{g}) ENERGY TRANSFER PROCESS

\underline{A. L. KALEDIN}, M. C. HEAVEN, and K. MOROKUMA, Cherry L. Emerson Center for Scientific Computation and Department of Chemistry Emory University, Atlanta, GA 30322.

Ab initio electronic structure calculations have been used to investigate the electronic energy transfer process:

I(^{2}P_{3/2})+O_{2}(a^{1}Δ_{g})↔I(^{2}P_{1/2})+O_{2}(X^{3}Σ^{−}_{g})

Potential energy surfaces for all states associated with the reactants and products were obtained using CASSCF and CASPT2 methods, including the effective one-electron spin-orbit Hamiltonian. Surfaces correlating with the reactants and products were all found to be non-bonding. Shallow van der Waals minima were predicted at long range. Surface crossings were found at energies below the I(^{2}P_{3/2})+O_{2}(a^{1}Δ_{g}) asymptote. It is probable that these crossings are responsible for the efficient transfer of electronic energy in this system.